OpenStudy (anonymous):

5[3(5y-4)-2(7y+6)]+3y-8=0?

OpenStudy (anonymous):

for this you have to use PEMDAS it means the order of how you solve things. it stands for Parantheses, Exponents, Multiply, Divide, Add, and then Subtract

OpenStudy (anonymous):

so, first you have to do what is inside the parantheses, and that is 3(5y-4)-2(7y+6)

OpenStudy (anonymous):

if you do distributive property, then that would be 15y-12-14y-12

OpenStudy (anonymous):

and then, you combine the like terms, so it equals to y-24

OpenStudy (anonymous):

so you then plug that into the equation which gives you 5(y-24) +3y-8=0

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

and then you do distributive property again which then converts it to 5y-120+3y-8=0

OpenStudy (anonymous):

and then if you combine it it equals to 2y+112=0

OpenStudy (anonymous):

then you subtract 112 on both sides, which gives you 2y=-112 then you divide 2 on both sides, which give you y=-56